# Signature-Informed Transformer for Asset Allocation

### v1

@@
Abstract
Robust asset allocation is a key challenge in quantitative finance, where deep-learning forecasters often fail due to objective mismatch and error amplification. We introduce the Signature-Informed Transformer (SIT), a novel framework that learns end-to-end allocation policies by directly optimizing a risk-aware financial objective. SIT’s core innovations include path signatures for a rich geometric representation of asset dynamics and a signature-augmented attention mechanism embedding financial inductive biases, like lead-lag effects, into the model. Evaluated on daily S&P 100 equity data, SIT decisively outperforms traditional and deep-learning baselines, especially when compared to predict-then-optimize models. These results indicate that portfolio-aware objectives and geometry-aware inductive biases are essential for risk-aware capital allocation in machine-learning systems. The code is available at: https://github.com/Yoontae6719/Signature-Informed-Transformer-For-Asset-Allocation
@@ 1 Introduction
Figure 1: A depiction of flawed deep learning strategies for asset allocation.
The predominant deep learning paradigm for this problem, illustrated in Figure 1, is a decoupled, two-stage pipeline: a forecasting model first predicts asset returns, and these predictions are then fed into a downstream portfolio optimizer (Moody & Saffell, 2001). This approach has drawbacks and suffers from two critical issues. First, the forecasting models typically employed are general-purpose architectures. They lack the financial inductive biases necessary to model the idiosyncratic structures of financial markets, such as the intricate lead-lag relationships between assets. Without a model architecture that explicitly reflects market dynamics, such models struggle to distinguish genuine signals from noise. Second, and more critically, this pipeline creates an objective mismatch that leads to error amplification. The forecaster is trained to minimize a statistical metric like the Mean Squared Error (MSE), i.e. the average squared difference between estimated and actual values. This objective is agnostic to the downstream task of portfolio construction, where even minuscule prediction errors can be magnified by the optimizer into volatile and impractical portfolio weights. Furthermore, an MSE objective implicitly incentivizes the model to favor assets that are easier to predict, potentially not considering harder-to-predict assets with larger estimation errors and distorting the final allocation. We argue that a robust solution requires moving beyond this fragile pipeline. The challenge is to develop a unified policy that learns an end-to-end mapping from market data directly to portfolio weights while being architecturally designed to model the known geometric properties of financial time series (Buehler et al., 2019; Hwang et al., 2025).
@@
    Path-wise Feature Representation: To better capture the complex dynamics of assets, the model generates features from each asset’s price history using Rough Path Signatures. This technique offers a principled summary of a path’s shape, encoding its trends and oscillations to provide a richer basis for decision-making (Lyons, 1998; Lyons & McLeod, 2022).
@@
2 Related works
Deep Learning in Asset Allocation
The application of deep learning in quantitative trading has largely bifurcated into two distinct paradigms. The first, the classic Predict Focused Learning (PFL) pipeline, focuses on developing return-prediction models. In this stream of research, complex architectures map market data to future price movements. For instance, Transformers have been adapted to capture temporal dependencies in asset prices for return forecasting (Fischer & Krauss, 2018; Yoo et al., 2021; Lim et al., 2021). Some models employ Graph Neural Networks (GNNs) to explicitly model inter-asset relationships, such as sector correlations, to improve prediction accuracy (Xu et al., 2021; Duan et al., 2025). Despite their architectural novelty, these methods inherit the fundamental flaws of a decoupled approach (Lee et al., 2024b). They suffer from objective mismatch, as optimizing for prediction error (e.g., Mean Squared Error) does not guarantee profitable portfolio construction, and are susceptible to error amplification, where small prediction inaccuracies lead to drastically suboptimal and unstable allocations (Chung et al., 2022). A more promising direction, which we term Decision Focused Learning (DFL), seeks to overcome these limitations by training policies end-to-end. These models learn a direct mapping from market state to portfolio allocations, optimizing a true financial objective like a risk-adjusted return metric. Foundational work demonstrated how to embed financial operators, such as portfolio value and Sharpe ratio, within a deep network, making the entire strategy differentiable and trainable via gradient descent (Buehler et al., 2019; Zhang et al., 2020; Costa & Iyengar, 2023). Recent research has increasingly emphasized embedding practical portfolio constraints into the model training phase. Typical examples include prohibiting short selling, ensuring full investment (i.e., portfolio weights sum to one), and placing upper or lower bounds on individual asset allocations, all of which are incorporated directly into the model architecture or loss function (Lee et al., 2024a; Hwang et al., 2025). While these end-to-end frameworks efficiently align the model’s training objective with financial goals, they often fall short in explicitly guiding the model to learn and utilize the diverse information present in multi-asset settings. This leaves a critical research gap. These models lack a strong financial inductive bias to explicitly represent the non-linear, path-dependent nature of price series and the geometric, time-local lead-lag relationships between assets. In our implementation, the predicted returns $\hat{\mu}$ serve only as internal logits for a differentiable allocation layer. All parameters are trained end-to-end solely through the portfolio-level CVaR objective, not a pointwise prediction loss, aligning with decision-focused learning. Our work addresses this gap by integrating the mathematical theory of path signatures directly into a transformer’s attention mechanism, creating an optimization-aware model that is architecturally designed to understand the underlying geometry of market dynamics. For general background on deep learning for portfolio optimization see (Zhang & Zohren, 2025; Lee et al., 2024b)
Transformer-Based Time Series Forecasting
The success of the Transformer architecture in natural language processing has inspired its widespread adoption for time series forecasting. The core innovation, the self-attention mechanism, allows these models to dynamically weigh the importance of all past time steps when predicting future values, enabling them to capture complex, long-range dependencies without the sequential processing limitations of recurrent neural networks(Vaswani et al., 2017; Li et al., 2019). To extend the receptive field without incurring the quadratic cost of full attention, a stream of variants introduce sparsity or hierarchical structure. For example, LogSparse(Li et al., 2019), ProbSparse(Zhou et al., 2021) and related kernels discard low-magnitude query–key interactions to achieve $\mathcal{O}(L\log L)$ complexity while retaining global context. From a more fundamental time series data perspective, Autoformer(Wu et al., 2021), FEDformer(Zhou et al., 2022) and ETSformer(Woo et al., 2022) decompose signals into trend–seasonality (or frequency-domain) components so that long-horizon patterns can be modeled additively and multiplicatively with reduced error accumulation. More recent PatchTST(Nie et al., 2022) and TimesNet(Wu et al., 2022) patch neighboring observations or convolve multi-scale windows before attention, embedding stronger inductive biases for periodicity and aliasing control. While these innovations alleviate the long-range dependency bottleneck, they remain largely data-agnostic. When applied to financial series they struggle with regime-dependent non-stationary, heavy-tailed noise, and asynchronous cross-asset lead–lag effects, causing attention scores to lock onto transient outliers and degrading out-of-sample robustness (Cartea et al., 2023; Cont, 2001; Miori & Cucuringu, 2022). Our approach departs from this paradigm by embedding each asset’s path in a Rough Path Signature space that is stable under time-reparameterization and robust to micro-structure noise, and by augmenting the attention logits with second-order cross-signature terms that encode the signed-area geometry underpinning lead–lag dynamics. Coupled with scenario-based optimization to hedge against structural breaks, SIT addresses both the generic long-range dependency problem and the finance-specific pathologies that limit existing Transformer forecasters.
Path Signatures in Time Series and Finance
The path signature, originating from Rough Path Theory, offers a non-parametric and faithful representation of streamed data by summarizing the geometry of a path as a sequence of iterated integrals (Lyons, 1998). A key property is its universality: any continuous function on the space of paths can be arbitrarily well-approximated by a linear function of the signature’s terms, making it a powerful basis for feature extraction (Chevyrev & Kormilitzin, 2016). In practice, the signature is truncated at a finite order $M$, yielding a vector $\mathrm{Sig}^{M}(\mathbf{X})$ that is robust to irregular sampling due to its invariance to time reparameterization. However, this truncation introduces a trade-off, as the feature dimension grows exponentially with the order $M$ and polynomially with the path dimension $d$, posing a significant computational burden. This challenge has motivated alternatives like signature kernels, which compute inner products in the high-dimensional feature space implicitly, avoiding explicit feature construction (Király & Oberhauser, 2019). In machine learning, signatures provide a potent inductive bias for modeling systems with path-dependent memory. The most direct application involves using truncated signatures as static input features for standard models (Gyurkó et al., 2013). More sophisticated integrations are found in continuous-time models like Neural Controlled Differential Equations (CDEs), which learn a vector field that is controlled by the input path, effectively modeling the system’s response to a driving signal (Kidger et al., 2020). For finance, a crucial insight arises from the signature’s geometry: the second-order terms of a joint signature over two asset paths precisely encode their signed area, a direct and robust measure of their temporal lead-lag relationship (Lyons & McLeod, 2022). This property has been successfully leveraged to build kernels for detecting asymmetric dependencies between financial instruments, offering a principled alternative to traditional correlation measures (Bonnier et al., 2019). Also, applications to finance span volatility/return modeling, derivatives, and market microstructure. Early studies extracted signature coordinates to forecast realized volatility and to detect temporal asymmetries (Gyurkó et al., 2013). In options, signatures parameterize no‑arbitrage dynamics and enable data‑driven pricing/hedging (Arribas et al., 2020), including transformer‑style encoders fed with log/signatures (Tong et al., 2023). A crucial geometric motif is the second‑order signed area,
$A(X^{i},X^{j})=\int X^{i}\,dX^{j}-\int X^{j}\,dX^{i},$ (1)
which encodes temporal asymmetry and lead–lag; signature kernels exploit this to compare pairs or small baskets of assets (Chevyrev & Kormilitzin, 2016; Király & Oberhauser, 2019). Our architecture operationalization this motif at scale: SIT injects cross‑asset signature information as a dynamic, query‑conditioned bias inside attention, so that pairwise signed‑area evidence modulates which assets attend to which others at each decision point (cf. Theorem 3.1). While signatures mitigate non‑stationarity and encode higher‑order interactions, they incur truncation bias and can suffer from a curse of dimensionality as either degree $M$ or the number of assets grows; kernelization trades feature savings for quadratic kernel costs (Salvi et al., 2021; Bonnier et al., 2019). Compared with state‑space or transformer baselines, signatures offer complementary bias—geometric invariances and lead–lag structure—rather than longer receptive fields alone. Prior signature‑based works typically (i) use signatures as fixed inputs or kernels outside the attention mechanism and (ii) optimize predictive losses, not portfolio objectives (Gyurkó et al., 2013; Tong et al., 2023; Bonnier et al., 2019). SIT differs by coupling signature‑augmented, cross‑asset attention with end‑to‑end CVaR optimization for long‑only, fully‑invested portfolios, aligning representation, interaction, and objective (Buehler et al., 2019).
@@ 3 Methodology
This section introduces the Signature-Informed Transformer (SIT), a novel approach to risk-aware portfolio allocation (Figure 2). After a brief overview of the problem and path signatures, we detail the model’s core components: (i) a unified embedding for signature, calendar, and asset features; (ii) a Signature-Informed Self-Attention mechanism that leverages cross-asset relations; and (iii) a CVaR-minimization training strategy for robustness to tail risk.
@@ 3.1 Preliminaries
A key ingredient in our framework is the use of path signatures to capture high-order variations and cross-asset interactions in price trajectories. For a continuous path $\mathbf{X}:[s,t]\to\mathbb{R}^{d}$, the signature $\text{Sig}(\mathbf{X}_{[s,t]})$ lies in the tensor algebra $\oplus_{k=0}^{\infty}(\mathbb{R}^{d})^{\otimes k}$. When truncated at level $M$, it becomes a finite-dimensional vector denoted as $\text{Sig}^{M}(\mathbf{X}_{[s,t]})=(1,\int_{s}^{t}d\mathbf{X}_{u},\int_{s}^{t}\int_{s}^{u}d\mathbf{X}_{r}\otimes d\mathbf{X}_{u},\ldots)$. In our financial context, $\mathbf{X}$ corresponds to the price process $\mathbf{S}_{t}$. First-order signature terms capture net increments for each asset, while second-order terms encode signed areas, revealing non-trivial correlations and lead-lag effects. For clarity, the key notations are provided in Appendix A.
Theorem 3.1 (Strict Lead-Lag Implies Positive Second-Order Signature).
Let $\mathbf{X}_{t}=(X^{1}_{t},X^{2}_{t})$ for $t\in[0,T]$ satisfy a strict lead-lag structure of Definition B.1. Then the second-level signature cross-term
@@
which yields $k$-step-ahead expected returns for $k=1,\ldots,K$. Portfolio weights for step $k$ are then obtained via $\mathbf{w}^{(k)}_{t_{i}}(\theta)=\text{softmax}\!(\hat{\boldsymbol{\mu}}^{(k)}_{t_{i}}(\theta)/\tau)\in\mathbb{R}^{d}$, ensuring a long-only, fully invested allocation at each future step. Note that $\hat{\boldsymbol{\mu}}^{1:K}_{t_{i}}$ is not trained with a prediction loss. It acts as the logits of the allocation layer, and gradients flow only from the portfolio objective below. Let $\mathbf{r}_{t_{i+k}}$ be the vector of realized asset returns over $t_{i+k-1},t_{i+k}]$, and define the corresponding portfolio loss $L^{(k)}_{t_{i+k}}(\theta_{\omega})=-(\mathbf{w}^{(k)}_{t_{i}}(\theta))^{\top}\mathbf{r}_{t_{i+k}}(\omega)$. The parameters $\theta$ are optimized by minimizing the expected Conditional Value-at-Risk (CVaR) of the $K$-step loss sequence within a scenario:
@@ 3.2 Signature-Informed Transformer (SIT)
This vector $\mathbf{q}^{\text{dyn}}_{k,j}$ represents the informational need of asset $j$ at slice $k$. The dynamic attention bias, $b_{k,h,j,l}$, for each head $h$, query asset $j$, and key asset $l$ at time slice $k$, is computed via an inner product:
@@
$z_{j,m}=\frac{(\mathbf{Q}_{k,h}\mathbf{K}_{k,h}^{\top})_{j,m}}{\sqrt{d_{k}}}+\gamma\langle(\mathbf{q}^{\mathrm{dyn}}_{k,j})_{h},(\boldsymbol{\beta}_{i,j,m})_{h}\rangle,\qquad\alpha_{j,m}=\frac{e^{z_{j,m}}}{\sum_{r=1}^{d}e^{z_{j,r}}},$ (13)
$D^{(\boldsymbol{\beta})}_{(\mathbf{q}^{\mathrm{dyn}}_{k,j})_{h}}\alpha_{j,l}=\gamma\,\alpha_{j,l}(1-\alpha_{j,l})\,\|(\mathbf{q}^{\mathrm{dyn}}_{k,j})_{h}\|_{2}^{2}>0.$ (14)
@@
Let $\mathbf{r}_{t_{i+k}}$ denote the realized asset-return vector over $t_{i+k-1},t_{i+k}]$. The step-$k$ portfolio loss is $L^{(k)}_{t_{i+k}}(\theta_{\omega})=-(\mathbf{w}^{(k)}_{t_{i}}(\theta))^{\top}\mathbf{r}_{t_{i+k}}(\omega)$. The overall objective is formally stated as:
@@
No auxiliary prediction losses are used. Eq. (18) is the sole training signal, avoiding the objective-mismatch issues discussed in Section  2. For each scenario $\omega$, the inner $\text{CVaR}_{\alpha}$ is taken over the intra-scenario empirical distribution. The following derivation shows the dual form and its empirical counterpart used for optimization:
@@ 4.1 Implementation details
Experiments used three portfolios of 30, 40, and 50 S&P 100 companies. The daily price data was sourced from Wharton Research Data Services (WRDS). The data was partitioned chronologically into distinct training, validation, and test periods. The training set spans from January 1, 2000, to December 31, 2016. The validation set from January 1, 2017, to December 31, 2019 and the test set from January 1, 2020, to December 27, 2024. This split covers multiple market regimes, including the recent volatility.
Baseline Models
The performance of our proposed model, SIT, is compared against a comprehensive suite of benchmarks spanning traditional and deep learning approaches. Traditional baselines include Equally Weighted Portfolio (EWP) (DeMiguel et al., 2009), Global Minimum Variance (GMV) (Clarke et al., 2011; Markowitz, 1952), Conditional Value-at-Risk (CVaR) (Rockafellar et al., 2000) and Hierarchical Risk Parity (HRP) (Lopez de Prado, 2016). The portfolio optimization strategy forms the second stage of our deep learning-based comparisons, which use predictions from various state-of-the-art time-series forecasting models as input. These forecasters include deep learning models such as Autoformer (Wu et al., 2021), DLinear (Zeng et al., 2023) FEDformer (Zhou et al., 2022), PatchTST (Nie et al., 2022), iTransformer (Liu et al., 2023), Non-stationary Transformers (NSformer) (Liu et al., 2022) and TimesNet (Wu et al., 2022). Details of the parameter search space are provided in Appendix  C.
Evaluation Metrics
4.2 Can SIT deliver superior risk-adjusted performance?
We evaluate the out-of-sample portfolio management efficacy of our proposed model: SIT. The comprehensive performance metrics, including risk-adjusted returns and downside risk, are presented for the 40- and 50-asset universes (see Appendix E for the 30-asset universe experiment). Our analysis underscores that the quality of asset allocation, rather than raw predictive accuracy, is the decisive factor for success, a central tenet of our work.
The empirical results, summarized in Table 1, demonstrate that SIT consistently and significantly outperforms all baseline models across the primary metrics of risk-adjusted return and wealth generation. In the 40-asset universe, for instance, SIT achieves a Sharpe Ratio of 0.6717 and a Sortino Ratio of 0.8232, decisively surpassing the next-best traditional baseline EWP and all deep learning counterparts. This translates into superior capital growth, with SIT yielding a Final Wealth Factor of 1.7903, the highest among all tested strategies.
Panel A. Asset 40 Universe
@@ 4.2 Can SIT deliver superior risk-adjusted performance?
Panel B. Asset 50 Universe
@@
The primary contribution of SIT becomes evident when contrasted with the predict-then-optimize models. These models, which rely on minimizing statistical forecasting error, exhibit poor and highly unstable portfolio performance. Many fail to outperform even simple heuristics. Their high standard deviations across runs underscore the problem of error amplification, where small prediction inaccuracies are magnified by the downstream optimizer into fragile, impractical allocations. This finding empirically validates our core hypothesis. Optimizing for prediction is not a valid proxy for optimizing for allocation quality. In addition to its inductive biases designed for financial assets, SIT’s decision-focused approach directly minimizes the portfolio’s CVaR, fundamentally aligning the model’s objective with the financial goal and thereby avoiding this critical pitfall.
Furthermore, SIT demonstrates a superior risk-return profile compared to traditional quantitative strategies. While risk-minimizing models like Global Minimum Variance (GMV) achieve low Maximum Drawdown (MDD) (e.g., 0.2743 in the 40-asset case), they do so at the cost of substantially lower returns (Sharpe Ratio of 0.4148). SIT, conversely, maintains a competitive MDD (0.3611) while delivering significantly higher returns.

### v2

@@
Signature-Informed Transformer for Asset Allocation
Yoontae Hwang    Stefan Zohren
Machine Learning, ICML, Asset Allocation, Trading
Modern deep learning for asset allocation typically separates forecasting from optimization. We argue this creates a fundamental mismatch where minimizing prediction errors fails to yield robust portfolios. We propose the Signature Informed Transformer to address this by unifying feature extraction and decision making into a single policy. Our model employs path signatures to encode complex path dependencies and introduces a specialized attention mechanism that targets geometric asset relationships. By directly minimizing the Conditional Value at Risk we ensure the training objective aligns with financial goals. We prove that our attention module rigorously amplifies signature derived signals. Experiments across diverse equity universes show our approach significantly outperforms both traditional strategies and advanced forecasting baselines. The code is available at: https://anonymous.4open.science/r/Signature-Informed-Transformer-For-Asset-Allocation-DB88
@@ 1 Introduction
Figure 1: A depiction of flawed strategies for asset allocation
The predominant deep learning paradigm for this problem, illustrated in Figure  1, is a decoupled, two-stage pipeline: a forecasting model first predicts asset returns, and these predictions are then fed into a downstream portfolio optimizer (Moody and Saffell, 2001). This approach has drawbacks and suffers from two critical issues. First, the forecasting models typically employed are general-purpose architectures. They lack the financial inductive biases necessary to model the idiosyncratic structures of financial markets, such as the intricate lead-lag relationships between assets. Without a model architecture that explicitly reflects market dynamics, such models struggle to distinguish genuine signals from noise. Second, and more critically, this pipeline creates an objective mismatch that leads to error amplification. The forecaster is trained to minimize a statistical metric like the Mean Squared Error (MSE), i.e. the average squared difference between estimated and actual values. This objective is agnostic to the downstream task of portfolio construction, where even minuscule prediction errors can be magnified by the optimizer into volatile and impractical portfolio weights. Furthermore, an MSE objective implicitly incentivizes the model to favor assets that are easier to predict, potentially not considering harder-to-predict assets with larger estimation errors and distorting the final allocation. We argue that a robust solution requires moving beyond this fragile pipeline. The challenge is to develop a unified policy that learns an end-to-end mapping from market data directly to portfolio weights while being architecturally designed to model the known geometric properties of financial time series (Buehler et al., 2019; Hwang et al., 2025a).
@@
    Path-wise Feature Representation: To better capture the complex dynamics of assets, the model generates features from each asset’s price history using Rough Path Signatures. This technique offers a principled summary of a path’s shape, encoding its trends and oscillations to provide a richer basis for decision-making (Lyons, 1998; Lyons and McLeod, 2022).
@@ 2 Methodology
This section introduces the Signature-Informed Transformer (SIT), a novel approach to risk-aware portfolio allocation (Figure 2). All relevant literature can be found in the Appendix A. After a brief overview of the problem and path signatures, we detail the model’s core components.
@@ 2.1 Preliminaries
A key ingredient in our framework is the use of path signatures to capture high-order variations and cross-asset interactions in price trajectories. For a continuous path $\mathbf{X}:[s,t]\to\mathbb{R}^{d}$, the signature $\text{Sig}(\mathbf{X}_{[s,t]})$ lies in the tensor algebra $\oplus_{k=0}^{\infty}(\mathbb{R}^{d})^{\otimes k}$. When truncated at level $M$, it becomes a finite-dimensional vector denoted as $\text{Sig}^{M}(\mathbf{X}_{[s,t]})=(1,\int_{s}^{t}d\mathbf{X}_{u},\int_{s}^{t}\int_{s}^{u}d\mathbf{X}_{r}\otimes d\mathbf{X}_{u},\ldots)$. In our financial context, $\mathbf{X}$ corresponds to the price process $\mathbf{S}_{t}$. First-order signature terms capture net increments for each asset, while second-order terms encode signed areas, revealing non-trivial correlations and lead-lag effects. For clarity, the key notations are provided in Appendix B.
Proposition 2.1 (Strict Lead-Lag Implies Positive Second-Order Signature (cf. (Chevyrev and Kormilitzin, 2016))).
Let $\mathbf{X}_{t}=(X^{1}_{t},X^{2}_{t})$ for $t\in[0,T]$ satisfy a strict lead-lag structure of Definition C.1. Then the second-level signature cross-term
@@
which yields $k$-step-ahead expected returns for $k=1,\ldots,K$. Portfolio weights for step $k$ are then obtained via $\mathbf{w}^{(k)}_{t_{i}}(\theta)=\text{softmax}\!(\hat{\boldsymbol{\mu}}^{(k)}_{t_{i}}(\theta)/\tau)\in\mathbb{R}^{d}$, ensuring a long-only, fully invested allocation at each future step. Note that $\hat{\boldsymbol{\mu}}^{1:K}_{t_{i}}$ is not trained with a prediction loss. It acts as the logits of the allocation layer, and gradients flow only from the portfolio objective below. Let $\mathbf{r}_{t_{i+k}}$ be the vector of realized asset returns over $t_{i+k-1},t_{i+k}]$, and define the corresponding portfolio loss $L^{(k)}_{t_{i+k}}(\theta_{\omega})=-(\mathbf{w}^{(k)}_{t_{i}}(\theta))^{\top}\mathbf{r}_{t_{i+k}}(\omega)$. The parameters $\theta$ are optimized by minimizing the expected Conditional Value-at-Risk (CVaR) of the $K$-step loss sequence within a scenario as
@@ 2.2 Signature-Informed Transformer (SIT)
This vector $\mathbf{q}^{\text{dyn}}_{k,j}$ represents the informational need of asset $j$ at slice $k$. The dynamic attention bias, $b_{k,h,j,l}$, for each head $h$, query asset $j$, and key asset $l$ at time slice $k$, is computed via an inner product as
@@
$z_{j,m}$ $=\frac{(\mathbf{Q}_{k,h}\mathbf{K}_{k,h}^{\top})_{j,m}}{\sqrt{d_{k}}}+\gamma\langle(\mathbf{q}^{\mathrm{dyn}}_{k,j})_{h},(\boldsymbol{\beta}_{i,j,m})_{h}\rangle$ (12)
$\alpha_{j,m}$ $=\frac{e^{z_{j,m}}}{\sum_{r=1}^{d}e^{z_{j,r}}},$ (13)
[…]
@@
Let $\mathbf{r}_{t_{i+k}}$ denote the realized asset-return vector over $t_{i+k-1},t_{i+k}]$. The step-$k$ portfolio loss is $L^{(k)}_{t_{i+k}}(\theta_{\omega})=-(\mathbf{w}^{(k)}_{t_{i}}(\theta))^{\top}\mathbf{r}_{t_{i+k}}(\omega)$. The overall objective is formally stated as
@@
No auxiliary prediction losses are used. Eq. (17) is the sole training signal, avoiding the objective-mismatch issues discussed in Section  A. For each scenario $\omega$, the inner $\text{CVaR}_{\alpha}$ is taken over the intra-scenario empirical distribution. The following derivation shows the dual form and its empirical counterpart used for optimization:
@@ 3.1 Implementation details
Experiments used three portfolios of 30, 40, and 50 S&P 100 companies. We also selected two additional portfolios of 10 and 20 assets from the DOW30 to validate performance against a different index composition, which is often characterized as more concentrated. Furthermore, to evaluate robustness across different market dynamics, we included two portfolios consisting of 50 and 100 assets from the CSI 300 index. The daily price data was sourced from Wharton Research Data Services (WRDS). The data was partitioned chronologically into distinct training, validation, and test periods. The training set spans from January 1, 2000, to December 31, 2016. The validation set from January 1, 2017, to December 31, 2019 and the test set from January 1, 2020, to December 27, 2024. This split covers multiple market regimes, including the recent volatility.
Panel A. Asset 40 Universe (S&P100)
@@
Panel B. Asset 50 Universe (S&P100)
@@
Baseline Models
The performance of our proposed model, SIT, is compared against a comprehensive suite of benchmarks spanning traditional and deep learning approaches. Traditional baselines include Equally Weighted Portfolio (EWP) (DeMiguel et al., 2009), Global Minimum Variance (GMV) (Clarke et al., 2011; Markowitz, 1952), Conditional Value-at-Risk (CVaR) (Rockafellar et al., 2000) and Hierarchical Risk Parity (HRP) (Lopez de Prado, 2016). The portfolio optimization strategy forms the second stage of our deep learning-based comparisons, which use predictions from various state-of-the-art time-series forecasting models as input. These forecasters include deep learning models such as Autoformer (Wu et al., 2021), DLinear (Zeng et al., 2023) FEDformer (Zhou et al., 2022), PatchTST (Nie et al., 2022), iTransformer (Liu et al., 2023), Non-stationary Transformers (NSformer) (Liu et al., 2022), TimesNet (Wu et al., 2022) and RFormer (Moreno-Pino et al., 2024). Details of the parameter search space are provided in Appendix  D.
Evaluation Metrics
3.2 Can SIT deliver superior risk-adjusted performance?
We evaluate the out-of-sample asset allocations efficacy of our proposed model: SIT. The comprehensive performance metrics, including risk-adjusted returns and downside risk, are presented for the 40- and 50-asset universes (see Appendix G for the 30-asset universe experiment). Our analysis underscores that the quality of asset allocation, rather than raw predictive accuracy, is the decisive factor for success, a central tenet of our work. The empirical results, summarized in Table 1, demonstrate that SIT consistently and significantly outperforms all baseline models across the primary metrics of risk-adjusted return and wealth generation. In the 40-asset universe, for instance, SIT achieves a Sharpe Ratio of 0.6717 and a Sortino Ratio of 0.8232, decisively surpassing the next-best traditional baseline EWP and all deep learning counterparts. This translates into superior capital growth, with SIT yielding a Final Wealth Factor of 1.7903, the highest among all tested strategies.
The primary contribution of SIT becomes evident when contrasted with the predict-then-optimize models. These models, which rely on minimizing statistical forecasting error, exhibit poor and highly unstable portfolio performance. Many fail to outperform even simple heuristics. Their high standard deviations across runs underscore the problem of error amplification, where small prediction inaccuracies are magnified by the downstream optimizer into fragile, impractical allocations. This finding empirically validates our core hypothesis. Optimizing for prediction is not a valid proxy for optimizing for allocation quality. In addition to its inductive biases designed for financial assets, SIT’s decision-focused approach directly minimizes the portfolio’s CVaR, fundamentally aligning the model’s objective with the financial goal and thereby avoiding this critical pitfall. Furthermore, SIT demonstrates a superior risk-return profile compared to traditional quantitative strategies. While risk-minimizing models like Global Minimum Variance (GMV) achieve low Maximum Drawdown (MDD) (e.g., 0.2743 in the 40-asset case), they do so at the cost of substantially lower returns (Sharpe Ratio of 0.4148). SIT, conversely, maintains a competitive MDD (0.3611) while delivering significantly higher returns. For the results on the DOW30 and SCI300, please refer to Appendix G.